# Watermarking Recommender Systems

### v1

@@ 4.2. Baseline and Model
One advantage of model watermarking is the ability to maintain the utility. We show it by comparing AOW with GRO (zhang2024defense, ), an defense method against model extraction attacks on recommender systems, which can also be used to protect the model intellectual property. GRO defines a swap loss and jointly trains with the original task. Although it can bring utility degradation to the attacker’s model, it will also harm the utility of the target model. We make an comparison between the utility of the target model under AOW and GRO, and demonstrate the effectiveness of AOW in preserving the utility. Following their work, we choose Bert4Rec (sun2019bert4rec, ) as the backbone model, and evaluate model performance by ranking all the items in the dataset. However, AOW is a model-agnostic approach that can be applied to arbitrary sequential models. We have also conducted experiments on another model SASRec (kang2018self, ), and the results and conclusions are similar with Bert4Rec. Therefore, we only present the results on Bert4Rec for simplicity and consistency.
@@ 4.3. Settings
We follow the hyperparameter setting from Bert4Rec for each dataset. For GRO, we follow their official implementation and setting. For distillation, we use the black-box model extraction attack proposed by Yue et al. (yue2021black, ) and follow their suggested setting to autoregressively generate 3,000 distillation sequences. For fine-tuning, we adopt the strategy by Yue et al. (yue2021black, ) and generate some new sequences autoregressively through querying the oracle model. The number of generated sequences vary between 1% to 20% of the number of sequences in the original dataset. These generated sequences serve as an approximation of the in-distribution data that an attacker might have access to, but are different from the data used to train the oracle model. We use them to fine-tune the watermarked model. We fine-tune each model for the same number of epochs as during training.

### v2

@@ 4.2. Baseline and Model
One advantage of model watermarking is the ability to maintain the utility. We show it by comparing AOW with GRO (Zhang et al., 2024), an defense method against model extraction attacks on recommender systems, which can also be used to protect the model intellectual property. GRO defines a swap loss and jointly trains with the original task. Although it can bring utility degradation to the attacker’s model, it will also harm the utility of the target model. We make an comparison between the utility of the target model under AOW and GRO, and demonstrate the effectiveness of AOW in preserving the utility. Following their work, we choose Bert4Rec (Sun et al., 2019) as the backbone model, and evaluate model performance by ranking all the items in the dataset. However, AOW is a model-agnostic approach that can be applied to arbitrary sequential models. We have also conducted experiments on another model SASRec (Kang and McAuley, 2018), and the results and conclusions are similar with Bert4Rec. Therefore, we only present the results on Bert4Rec for simplicity and consistency.
@@ 4.3. Settings
We follow the hyperparameter setting from Bert4Rec for each dataset. For GRO, we follow their official implementation and setting. For distillation, we use the black-box model extraction attack proposed by Yue et al. (Yue et al., 2021a) and follow their suggested setting to autoregressively generate 3,000 distillation sequences. For fine-tuning, we adopt the strategy by Yue et al. (Yue et al., 2021a) and generate some new sequences autoregressively through querying the oracle model. The number of generated sequences vary between 1% to 20% of the number of sequences in the original dataset. These generated sequences serve as an approximation of the in-distribution data that an attacker might have access to, but are different from the data used to train the oracle model. We use them to fine-tune the watermarked model. We fine-tune each model for the same number of epochs as during training.